# Problem Solved? Information Extraction Design Space for Layout-Rich Documents using LLMs

### v1

@@ 4.2 The Input Dimension
We substitute OCR input with Markdown and as outcomes in both STL and UTL scenarios. The differences in performance between OCR and Markdown are model- and context-dependent, exhibiting no consistent trend favoring one input type over the other, as shown in Table 2.
@@
OCR input serves as a stable baseline for IE tasks, delivering consistent performance across models. GPT-4o has noticeable performance drops with Markdown input, indicating its reliance on OCR for optimal results. In contrast, Markdown marginally improves performance for LLaMA3-70B at both STL and UTL scenarios, suggesting its potential benefits from the additional structure or semantic cues. GPT-3.5 demonstrates robustness to changes in input type, with only slight fluctuations in performance. On average, OCR marginally outperforms Markdown (0.650 vs. 0.648), but the differences are minor, with standard deviations indicating similar stability.
@@ 4.3 The Chunk Dimension
To evaluate the impact of chunk size, we varied it from medium to max and small while keeping all other parameters constant. Table 3 demonstrates how chunk size affects performance across STL and UTL levels.
@@
Medium and max chunk sizes provide the most consistent and stable results across models, with an average F1 score of 0.650 (±0.011) and 0.651 (±0.011), respectively. Due to insufficient context, small chunk sizes result in significant performance drops, particularly for GPT-3.5 and GPT-4o. These findings suggest that max chunk size is optimal, but medium can be a good option for LLMs with limited context lengths..
@@ 4.4 The Prompt Dimension
Table 4 presents the impact of prompt type and the number of examples on model performance at STL and UTL levels. Surprisingly, in-context demonstrations do not enhance performance for either few-shot or CoT experiments. For both experiments, the setting with zero examples achieves the highest average performance: few-shot 0.650 (±0.011) and CoT 0.649 (±0.008). Performance consistently declines as the number of examples increase, likely due to noise that impairs generalization. Overall, there is no significant difference between few-shot and CoT.
@@ 4.5 Output Refinement
Schema Mapping
involves mapping the predicted schema keys to the target schema keys. Our results show no change in F1 scores compared to the initial predictions. This suggests that the models already effectively return the correct attributes, making the mapping step unnecessary.
Data Cleaning
uses the defined data types to perform automatic value cleaning, consistently achieving the highest F1 scores across all models. This underscores the need for post-processing steps for IE with LLMs to align the extracted data with the target format to handle LLM hallucinations and inconsistent source data formats (see Sec. 3.3).
@@ 4.6 Evaluation Techniques
We explore three evaluation techniques to assess their impact on model performance (see Sec. 3.3). On average, Fuzzy Match achieved the highest F1 score (0.733), outperforming Substring Match (0.676) and Exact Match, as shown in Table 6. We provide a detailed error analysis of fuzzy and substring match accuracy in Appendix A.3, showing that they provide a near-perfect precision when manually checked for semantic equivalence with precision scores of 0.98 and 1.00, respectively. This shows Fuzzy Match’s ability to balance flexibility and precision.
@@ 4.7 Putting it All Together
In the preceding sections, we investigated the influence of various parameters on model performance along the IE extraction pipeline, analyzing one factor at a time. Drawing from the underlying 12 experiments, we identified the optimal parameter for each step and each model based on the experimental outcomes (Table 7). In addition, we conducted an exhaustive full factorial exploration with 432 configurations ($2*3*2*4*3*3=432$, see Table 1) to find the best parametrization per LLM (Table 8). Lastly, we find the worst configuration on a per LLM and per model basis (Table 9). The performance of these different configurations is depicted in Figure 3. We gain several insights:
  • OFAT approximates well the Brute-Force configuration with a fraction ($\sim 2.8\%$) of the required computation. We see in Table 7 and Table 8 that they match except for 4 parameter choices (prompt and example No. parameters for GPT-4o and LlaMa3). Likewise, their F1 scores are close to each other (Figure 3), with OFAT achieving 0.791 and Brute-Force achieving 0.801 overall.
  • Adapting the IE pipeline to the LLM is necessary to achieve a competitive performance. Overall, the OFAT configuration improves from the baseline F1 of 0.650 to 0.791, a $22\%$ improvement. For Brute-Force, the improvement is $23\%$. In comparison, the worst configuration achieves only 0.45 on average, almost half of the best configuration.
  • Besides the need for pipeline customization, common patterns exist across LLMs. First, output refinement and evaluation techniques boost performance across all models. Second, there is a tendency to larger context sizes. Lastly, larger models (GPT-4o, LLaMa3) benefit more from examples, while the CoT pattern generally aids IE.
@@
Chunk Size Medium Max Max
Prompt CoT CoT CoT
@@
Table 9: Worst performance configurations on a per-model basis and corresponding performance results.
Parameter GPT-3.5 GPT-4o LLaMA3
Input Type OCR Markdown OCR
Chunk Size Small Small Small
Prompt Few-shot Few-shot Few-shot
Example No. 5 5 5
Output Refin. Initial Initial Initial
Evaluation Exact Exact Exact
Figure 3: F1 scores of different LLMs across three configurations (Baseline, OFAT, and Brute-Force). Each bar represents the mean F1 score of a model for the corresponding configuration.
Finally, we compare our purely text-based approach to IE to (1) GPT4-vision, a multimodal LLM, and (2) LayoutLMv3, a leading layout-aware model that we fine-tune to our dataset (Table 10). Despite fine-tuning, LayoutLMv3 does not perform as well as our evaluated LLMs, even on documents that contain the same structure as the fine-tuning data (STL). We also observe that LLMs remain stable across template types. The best-performing model is GPT4-vision, a multi-modal LLM, where we directly provide an image of the PDF for IE. The model benefits from less context loss due to chunking and its ability to jointly use textual and visual features. However, when considering token usage and cost, GPT4-vision requires $\sim 2$ times the tokens compared to text-only approaches and $>10$ times the cost under the current OpenAI pricing scheme (Nov. 2024). Therefore, text-only approaches constitute a good trade-off between performance and cost for practical applications.
Table 10: Cost (per experiment) and performance of LLMs across STL and UTL levels for the OFAT configuration. The tokens for LayoutLMv3 include the fine-tuning. We run LauoutLMv3 and LlaMa3 locally, without API costs.
@@
GPT-3.5 322K \$0.18 0.778 0.754
GPT-4o 278K \$0.40 0.797 0.790
LLaMA3 239K \$0 0.843 0.820
GPT-4-vision 585K \$4.76 0.902 0.897
@@ 5 Related Work
Transformer-based models like BERT Devlin et al. (2019), GPT Brown et al. (2020), and RoBERTa Liu et al. (2019b) have advanced NLP with self-attention and large-scale pretraining but struggle with layout-rich documents (LRDs). To address this, layout-aware and multimodal models have emerged. LayoutLM Xu et al. (2020b) integrates spatial features, with later versions Xu et al. (2021, 2022) and models like GPT-4V OpenAI (2023) and Gemini Pro Anil et al. (2023) enhancing document understanding through multimodal learning. LayoutLLM Fujitake (2024) and structural-aware approaches Lee et al. (2022) further improve extraction accuracy, while end-to-end models like Donut Kim et al. (2022) bypass OCR for direct document image processing. Additionally, ChatUIE Xu et al. (2024a) adopts a chat-based approach for flexible IE, while ReLayout Jiang et al. (2025) enhances document understanding through layout-aware pretraining, advancing LLMs for real-world use.
Strategies for IE from LRDs.
@@
Graph-based models like GCNs Liu et al. (2019a) and AliGATr Nourbakhsh et al. (2024) enhance relation extraction by capturing textual-visual relationships. Reading order is critical; Token Path Prediction (TPP) Zhang et al. (2023) resolves OCR layout ambiguities, while global tagging Shaojie et al. (2023) mitigates text ordering issues for better extraction. For structured data, TabbyPDF Jain et al. (2020) targets table extraction, DocExtractor Zhong et al. (2020) processes forms, and LMDX Perot et al. (2024) unifies OCR, preprocessing, and postprocessing for document IE. Additionally, XFORMPARSER Cheng et al. (2025) offers a simple yet effective multimodal and multilingual approach for parsing semi-structured forms.
Preprocessing, Chunking, Prompting, Postprocessing, and Evaluation Techniques.
@@
Chain-of-Thought (CoT) prompting Wei et al. (2022) enhances reasoning in complex LRD extraction tasks, while diverse prompt-response datasets Zmigrod et al. (2024a) improve LLM robustness. Instruction-finetuned LLMs have also demonstrated effectiveness in domain-specific applications, such as clinical and biomedical tasks, where zero-shot and few-shot learning enable adaptive extraction without extensive fine-tuning Labrak et al. (2024). Postprocessing techniques, including text normalization, entity resolution Hwang et al. (2021), and majority voting Wang et al. (2022a), refine extracted data by correcting OCR and extraction errors for greater accuracy.
Despite advancements, current studies often focus on isolated components rather than evaluating full IE pipelines. Key stages such as OCR, chunking, and postprocessing are frequently assessed independently, leading to an incomplete understanding of their interplay. Our research bridges this gap by systematically evaluating strategies across all pipeline stages to identify optimal configurations for leveraging pre-trained models.
@@ 6 Conclusion
In this paper we introduce the design space for IE from LRDs using LLMs, with challenges of data structuring, model engagement, and output refinement. We consider the design choices within those challenges. We examine the effects of these choices on the overall performance through OFAT and full-factorial exploration of the parameters. Finally, we show that a well-designed general purpose LLM can provide a cost-effective alternative to the the specialized models for IE from LRDs.
@@
For the experimental evaluation in the design space, we focused on the Visually Rich Document Understanding (VRDU) dataset Wang et al. (2023b), which has ground-truth annotations that make it easy for result analysis. We believe incrementally adding new datasets focusing on layout-rich documents with ground truth and making a combined benchmark would help improve the design space itself as we can test different application-specific scenarios.
@@
The evaluation metrics are planned to be extended with metrics that are popularly used in the context of LLMs such as ROUGE Lin (2004) and BLEU Papineni et al. (2002) scores and their adaptations Yang et al. (2018). A fair comparison using cost-related metrics is mostly harder to compute due to their changing depending on the scenario. For instance, in some scenarios where we listed cost as 0 due to having a free-to-use service, it actually comes with a cost. Similarly, the cost of computation can be considered from different perspectives such as the energy usage of a model.
@@
The test suite is currently limited to the steps that we listed in Fig. 2, whereas one could imagine additional steps or factors which affect the performances and may even deliver more satisfactory outcomes. We would like to incorporate additional steps that we learn from the community and incrementally enlarge the design space and extend the testing capabilities for IE from LRDs.
Last, we have only evaluated three LLMs, one multi-modal LLM and one layout-aware, fine-tuned model (LayoutLMv3). In the future, we want to extend our evaluation to further LLMs and fine-tuned, layout-aware models (e.g., ). Despite that, our results already show a clear trend in which specialized, fine-tuned models struggle to keep up with progress in general-purpose LLMs. E.g., LayoutLMv3, only reaches an F1 score of 0.681 with fine-tuning and on the same document types (STL), substantially less than Llama3, the best text-based LLM with 0.832 and GPT-4-vision, our tested multi-modal LLM with 0.890. This underlines the importance of our work, in which choices in data structuring, model engagement, and output refinement become critical in achieving the best performance.

### v2

@@ 4.2 The Input Dimension
We substitute OCR input with Markdown and evaluate the resulting performance in both STL and UTL scenarios. The performance differences between OCR and Markdown inputs vary by model and context, showing no consistent trend favoring either input type, as shown in Table 2.
@@
OCR input serves as a stable baseline for IE tasks, delivering consistent performance across models. GPT-4o has noticeable performance drops with Markdown input, indicating its reliance on OCR for optimal results. In contrast, Markdown marginally improves performance for LLaMA3-70B at both STL and UTL scenarios, suggesting its potential benefits from the additional structure or semantic cues. GPT-3.5 demonstrates robustness to changes in input type, with only slight fluctuations in performance. On average, OCR marginally outperforms Markdown (0.650 vs. 0.648), but the differences are minor, with standard deviations indicating similar stability (see Appendix A.1 for detailed insights).
@@ 4.3 The Chunk Dimension
To evaluate the impact of chunk size, we vary it while keeping all other parameters constant. Table 3 demonstrates how chunk size affects performance across STL and UTL levels.
Medium and max chunk sizes provide the most consistent and stable results across models, with an average F1 score of 0.650 (±0.011) and 0.651 (±0.011), respectively. Due to insufficient context, small chunk sizes result in significant performance drops, particularly for GPT-3.5 and GPT-4o. These findings suggest that max chunk size is optimal, but medium can be a good option for LLMs with limited context lengths.
@@ 4.4 The Prompt Dimension
Table 4 shows the impact of prompt type and number of examples on model performance for STL and UTL. Surprisingly, in-context demonstrations do not enhance performance for few-shot or CoT experiments (see Appendix A.1 for detailed insights). For both experiments, the setting with zero examples achieves the highest average performance: few-shot 0.650 (±0.011) and CoT 0.649 (±0.008). Performance consistently declines as the number of examples increase, likely due to noise that impairs generalization. Overall, there is no significant difference between few-shot and CoT.
@@ 4.5 Output Refinement
Schema Mapping involves mapping the predicted schema keys to the target schema keys. Our results show no change in F1 scores compared to the initial predictions. This suggests that the models already effectively return the correct attributes, making the mapping step unnecessary.
Data Cleaning uses the defined data types to perform automatic value cleaning, consistently achieving the highest F1 scores across all models. This underscores the need for post-processing steps for IE with LLMs to align the extracted data with the target format to handle LLM hallucinations and inconsistent source data formats (see Sec. 3.3).
@@ 4.6 Evaluation Techniques
We explore three evaluation techniques to assess their impact on model performance (see Sec. 3.3). On average, Fuzzy Match achieved the highest F1 score (0.733), outperforming Substring Match (0.676) and Exact Match, as shown in Table 6.
To validate the reliability of these metrics, we manually analyzed cases where Exact Match failed but Substring or Fuzzy Match succeeded. This review confirmed that most soft-matched predictions were semantically correct, with errors largely due to OCR noise, formatting differences, or annotation inconsistencies. Notably, fuzzy match (with a strict 0.8 threshold) captured valid predictions that exact and substring match missed, such as truncated organization names. Our detailed error analysis in Appendix A.4 shows that fuzzy and substring matching provide near-perfect precision when manually checked for semantic equivalence, with precision scores of 0.98 and 1.00, respectively. This shows Fuzzy Match’s ability to balance flexibility and precision.
@@ 4.7 Putting it All Together
We investigated the influence of various parameters on model performance along the IE extraction pipeline, analyzing one factor at a time. Drawing from the underlying 12 experiments, we identified the optimal parameter for each step and each model based on the experimental outcomes (Table 7). To validate this approach, we conducted an exhaustive full factorial exploration with 432 configurations to identify the overall best-performing (Table 8) and worst-performing parameter combinations per LLM. For the comparative analysis in Figure 3, we use a single global OFAT configuration that achieves the highest average performance across all models (selection criteria in Appendix A.1). The performance of these different configurations, including the worst, is depicted in Figure 3. We gain several insights:
  • OFAT approximates well the best-performing Brute-Force configuration with a fraction ($\sim 2.8\%$) of the required computation. We see in Table 7 and Table 8 that they match except for 8 parameter choices (chunk size, prompt and example No. parameters). Likewise, their F1 scores are close to each other (Figure 3), with OFAT achieving 0.783 and Brute-Force achieving 0.801 overall.
  • Adapting the IE pipeline to the LLM is crucial for competitive performance. Overall, the OFAT configuration improves from baseline F1 of 0.650 to 0.783, a $20\%$ improvement. For the best-performing Brute-Force, the improvement is $23\%$. In comparison, the worst configuration achieves an average score of only 0.5, which is approximately 64.1% of the best configuration’s performance.
  • Common patterns across LLMs. First, output refinement and evaluation techniques boost performance for all LLMs. Second, there is a trend towards larger context sizes. Lastly, larger models (GPT-4o, LLaMa3) benefit more from examples, while the CoT pattern generally aids IE.
FUNSD Dataset. We provide the equivalent OFAT analysis for the FUNSD dataset in Appendix A.5. While we observed different patterns with respect to the IE pipeline configuration, our experiments on FUNSD also show a clear improvement through output refinement techniques and the use of tolerant evaluation metrics. Most importantly, they show that our OFAT method nearly reaches the best-performing Brute-Force performance (only $0.8$ F1 points lower), and improves 90% from the baseline IE results. This indicates that our OFAT method for IE pipeline configuration is generalizable and can help practitioners and researchers to find the best configuration for their specific dataset.
@@
Table 8: Best-performing Brute-Force configurations on a per-model basis and corresponding performance results.
@@
Chunk Size Max Medium Small
Prompt Few-shot CoT CoT
@@
Figure 3: F1 scores of different LLMs across three configurations (Baseline, OFAT, and Brute-Force). Each bar represents the mean F1 score of a model for the corresponding configuration. The OFAT configuration uses a single global parameter set optimized across all models.
Finally, we compare our text-based approach to IE to (1) GPT-4o-vision and Qwen2.5-vision, multimodal LLMs, (2) LayoutLMv3 and ERNIE-Layout, layout-aware models that we fine-tune to our dataset (Table 9). Despite fine-tuning, LayoutLMv3 and ERNIE-Layout perform worse than our evaluated LLMs, even on documents that contain the same structure as the fine-tuning data (STL). We also observe that LLMs remain stable across template types. The best-performing models are GPT-4o-vision followed by Qwen2.5-vision, both multimodal LLMs, where we directly provide an image of the PDF for IE. These models benefits from less context loss due to chunking and their ability to use textual and visual features jointly. However, when considering token usage and cost, GPT-4o-vision requires $\sim 2\times$ the tokens compared to text-only approaches and $10\times$ the cost under the current OpenAI pricing scheme (Nov. 2024), while Qwen2.5-vision uses $\sim 1.4\times$ the tokens with no API costs when run locally. Therefore, text-only approaches constitute a good trade-off between performance and cost for practical applications, with Qwen2.5-vision offering a viable middle ground.
Table 9: Cost (per experiment) and performance of LLMs across STL and UTL levels for the OFAT configuration. The tokens for LayoutLMv3 and ERNIE-Layout include the fine-tuning. We run Qwen2.5-vision, LayoutLMv3, ERNIE-Layout, and LlaMa3 locally, without API costs.
@@
GPT-3.5 322K \$0.18 0.777 0.776
GPT-4o 278K \$0.40 0.798 0.788
LLaMA3 239K \$0 0.784 0.776
GPT-4o-vision 585K \$4.76 0.902 0.897
Qwen2.5-vision 385K \$0 0.890 0.891
@@ 5 Related Work
IE using LLMs. Transformer-based models Devlin et al. (2019); Brown et al. (2020); Liu et al. (2019b) have advanced NLP with self-attention and large-scale pretraining but struggle with layout-rich documents (LRDs). To address this, layout-aware and multimodal models have emerged Xu et al. (2020b, 2021, 2022), including generative approaches Powalski et al. (2021). Additionally, some models enhance document understanding through multimodal learning OpenAI (2023); Anil et al. (2023). Structural-aware approaches Fujitake (2024); Lee et al. (2022) further improve extraction accuracy, while end-to-end models Kim et al. (2022) bypass OCR for direct document image processing. Furthermore, some approaches adopt a chat-based paradigm for flexible IE Xu et al. (2024a), while others Jiang et al. (2025) enhance document understanding through layout-aware pretraining, advancing LLMs for practical applications.
@@
Strategies for IE from LRDs. Graph-based models Liu et al. (2019a); Nourbakhsh et al. (2024) enhance relation extraction by capturing textual-visual relationships. A critical factor in layout-rich document (LRD) processing is maintaining the reading order, where Token Path Prediction (TPP)Zhang et al. (2023) resolves OCR layout ambiguities, and global taggingShaojie et al. (2023) mitigates text ordering issues, thereby improving extraction accuracy. For structured data, specific approaches target various document types: table extraction Jain et al. (2020), form processing Zhong et al. (2020), and unified OCR, preprocessing, and postprocessing for document IE Perot et al. (2024). Additionally, a simple yet effective multimodal and multilingual method for parsing semi-structured forms is presented in Cheng et al. (2025). Despite the variety of models proposed, comparative studies evaluating text-only, multimodal, and layout-aware LLMs remain limited Van Landeghem et al. (2023); Stanisławek et al. (2021).
@@
Preprocessing, Chunking, Prompting, Postprocessing, and Evaluation Techniques. Chain-of-Thought (CoT) prompting Wei et al. (2022) enhances reasoning in complex LRD extraction tasks, while diverse prompt-response datasets Zmigrod et al. (2024a) improve LLM robustness. Instruction-finetuned LLMs have also demonstrated effectiveness in domain-specific applications, such as clinical and biomedical tasks, where zero-shot and few-shot enable adaptive extraction without extensive fine-tuning Labrak et al. (2024). Postprocessing techniques, including text normalization, schema alignment Fürst et al. (2023), entity resolution Hwang et al. (2021), and majority voting Wang et al. (2022a), refine extracted data by correcting OCR and extraction errors.
Despite advances, current studies often focus on isolated components rather than evaluating complete IE pipelines, leading to an incomplete understanding of their interplay. Our findings, quantified and contextualized within a controlled and reproducible framework. Our study enables structured evaluation of IE techniques and systematically analyzes the trade-offs between performance and efficiency across diverse model categories.
@@ 6 Conclusion
This paper introduces the design space for IE from LRDs using LLMs, with challenges of data structuring, model engagement, and output refinement, considering the design choices within those challenges. We examine the effects of these choices on the overall performance through one-factor-at-a-time (OFAT) and full-factorial (Brute-Force) exploration of the parameters. We show that well-configured general-purpose LLMs provide an effective alternative to specialized models for IE from LRDs. Our cost-effective OFAT method (2.8% of required computation) generalizes across two datasets (VRDU and FUNSD) with a performance that is just 0.8–1.8% lower than the best-possible configuration that can be found through a full-factorial exploration. This makes our results and method directly applicable to IE practitioners who need to configure their IE pipeline for their dataset. Towards this goal, we have open-sourced LayIE-LLM, our layout-aware IE test suite (https://github.com/gayecolakoglu/LayIE-LLM).
@@
Our experimental evaluation focused on the Visually Rich Document Understanding (VRDU) dataset Wang et al. (2023c) and the FUNSD dataset Jaume et al. (2019). VRDU was chosen for its comprehensive OCR outputs, well-defined schema, and diverse template types (Amendment, Short-Form, and Dissemination) that capture complex layouts. It also supports generalization tasks–Single Template Learning (STL), Unseen Template Learning (UTL), and Mixed Template Learning (MTL)–making it suitable for visually rich document understanding. In contrast, FUNSD was selected due to its widespread use in document understanding tasks, particularly for extracting structured information from simpler, form-based layouts. While FUNSD does not share the layout diversity and complexity of VRDU, it serves as a useful baseline for form-centric document analysis. This provides a different perspective by offering insights into scenarios where key-value pairs are organized within relatively simple templates.
One key limitation arises from the differences in schema structure between these datasets. The VRDU dataset has structured templates that facilitate consistent key mapping during post-processing (see Section 3.3), while the FUNSD dataset features dynamic, document-specific key-value pairs, making mapping more challenging. To address this, we plan to enhance our existing schema-mapping step by incorporating context-aware strategies that dynamically adapt to variations and leverage semantic similarity to reduce mismatches.
Another limitation lies in the construction of few-shot prompts (see Appendix A.3). Currently, the LLM generates these prompts by condensing the original OCR data while preserving only text features, aiming to reduce token usage and cost. However, in the tested document, both text and layout features are preserved without modification, embedding spatial structure into text sequences to enhance model understanding. Improving the LLM-generated few-shot prompt process by better integrating spatial features directly into the prompts could enhance the model’s performance in complex scenarios.
@@
Our manual investigation (as detailed in Appendix A.1) demonstrated that fuzzy match is more tolerant of valid variations while maintaining high accuracy, as verified through manual evaluation; however, we acknowledge that it may relax correctness criteria and inflate performance metrics. To address this, we plan to extend the evaluation with metrics commonly used in the context of LLMs, such as ROUGE Lin (2004) and BLEU Papineni et al. (2002) scores and their adaptations Yang et al. (2018). A fair comparison using cost-related metrics is mostly harder to compute due to their changing depending on the scenario. For instance, for Llama3, ERNIE-Layout, and LayoutLMv3, we listed the cost as 0 due to having a free-to-use service; it comes with a cost. Similarly, the cost of computation can be considered from different perspectives, such as the energy usage of a model.
@@
The test suite is currently limited to the steps that we listed in Fig. 2, whereas one could imagine additional steps or factors that affect the performance and may even deliver more satisfactory outcomes. We would like to incorporate additional steps that we learn from the community and incrementally enlarge the design space, and extend the testing capabilities for IE from LRDs.
Another limitation stems from the OFAT strategy used to explore the parameter space. While OFAT enables systematic tuning with minimal computational cost, it inherently fails to detect interaction effects between parameter combinations. To evaluate whether this limitation affected our results, we performed a brute-force search as a benchmark, directly comparing the OFAT outcomes to the best results from the brute-force approach. The comparison showed that OFAT approximated the full-grid results well, indicating that, in our specific scenario, OFAT was sufficient despite its theoretical limitations. We consider this trade-off acceptable for our goal of providing a fast, reproducible tuning method for real-world LLM deployment. Nevertheless, future work may incorporate interaction-aware designs to further uncover parameter synergies.
Last limitation of our study is the variability in performance between the different types of models we tested when handling LRDs. While some models, such as LayoutLMv3 and ERNIE-Layout, are fine-tuned for layout-aware tasks, others, like LLaMA3 and GPT-4o-vision, are optimized for text processing or multi-modal inputs, with their relative strengths and weaknesses varying across the evaluated dimensions. This underscores the need for broader benchmarking with additional models, such as DocLLM, LayoutLLM, and ReLayout, to better understand performance differences.